OpenStudy (anonymous):

I need to write sin^2(x)+sin^2(x)tan^2(x) in factored form including only one function

OpenStudy (anonymous):

sin^2(x)[1 + tan^2(x)]

OpenStudy (anonymous):

The answer in the back of my Trig book says tan^2(x) is the correct answer, I just need to know how the book got that to show my work.

OpenStudy (anonymous):

sin^2(x)(1+tan^2(x))=sin^2(x)sec^2(x)=sin^2x/cos^2x=tan^2x